OpenStudy (anonymous):

Show that the taylor series works for N=1

OpenStudy (anonymous):

???????????

OpenStudy (goformit100):

First of all do you know Taylor Series ?

OpenStudy (anonymous):

yes

OpenStudy (goformit100):

Please write it

OpenStudy (anonymous):

\[\sum_{k=0}^{N} \frac{ (x-c)^{k} }{ k! } \times f ^{(k)}(c) + \int\limits_{c}^{x} \frac{ (x-t)^{N} }{ N! } \times f ^{(N+1)} (t) = f(x) \]